OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -11 0 -1 1 9 g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

@Thesmarterone or anyone pls

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@chlobohoe

OpenStudy (anonymous):

@Littlebird Plz Help Me

OpenStudy (anonymous):

@Littlebird

OpenStudy (anonymous):

What about part b @Littlebird

OpenStudy (littlebird):

Forgot a minus sign, so I reposted it

OpenStudy (littlebird):

f(x) = 10x + 1 g(x) = 5x + 1 y = mx + b (m is slope. b is y-intercept.)

OpenStudy (anonymous):

Alrght thx

OpenStudy (littlebird):

So in f(x) = 10x + 1... The slope is 10 and The y intercept is 1

OpenStudy (anonymous):

Is that for Part B @littlebird

OpenStudy (littlebird):

You can basically find the answer for both questions at once In part a the slope for f(x), which is 10, is larger than the slope for g(x), which is 5. The y intercepts are the same

OpenStudy (littlebird):

In these steps in the picture, I found both the slope and the y intercept of f(x) g(x) was already given to you